OpenStudy (anonymous):

Solve the following system of equations algebraically:? 20x+Y=400 10x+y=600.....

OpenStudy (anonymous):

Multiply the second equation by -2 and it will cancel out the x term. Then you have: 20x + y = 400 -20x - 2y = -1200 This gives us the new set of equations: -y = -800 y = 800 Plug in and solve for x. Hope this helps. Please like Tutor Sean on facebook, and visit my website at http://www.tutorsean.net If this answer helps, please choose Good Answer.

OpenStudy (anonymous):

why do u Multiply it by -2?

OpenStudy (anonymous):

where did u get the -2 from?

OpenStudy (anonymous):

The way that this works is if we multiply an entire equation by a constant, it does not change the equation. Kind of like 1 is the same as 2/2 which is the same as 4/4. Does this help?

OpenStudy (anonymous):

This is what we call the elimination method.

OpenStudy (anonymous):

well i'm still alittle lost.

OpenStudy (anonymous):

help?

OpenStudy (anonymous):

Sure. I have a video on youtube that explains the whole process. My youtube screen name is TutorSean. I don't have a video on it right now, but will soon. Let me continue to try to clarify for you.

OpenStudy (anonymous):

If we have two equations. x + 2y = 1 x + 3y = 4 Then we can say leave equation 1 alone, and then multiply equation 2 by -1. Then we have: x + 2y = 1 -x - 3y = -4 If we rewrite the two equations together we have: 2y = 1 -3y = -4 Then we add together. -y = -5 So y = 5 Then replacing y with negative 5 in either of the original equations gives us: x + 2(5) = 1 which means that x + 10 = 1, x = -9 Do it in the second equation: -x - 3(5) = -4 -x - 15 = -4 -x = 9 So x = -9 This shows that the system is what we call consistent. If you need some more help, please contact me through my webpage, http://www.tutorsean.net and I will be more than happy to help you through EMAIL or phone at no charge. Sean